OpenStudy (anonymous):

What are the missing values in the following geometric sequence? __, 8, 2, __, __.

OpenStudy (anonymous):

is it 14,8,2,-4,-10?

OpenStudy (texaschic101):

32, 8, 2, 1/2, 1/8 32 * 1/4 = 32/4 = 8 8 * 1/4 = 8/4 = 2 2 * 1/4 = 1/2 1/2 * 1/4 = 1/8

OpenStudy (texaschic101):

common ratio is 1/4

OpenStudy (anonymous):

OHHH Thank you!